# Evidence Contextualization and Counterfactual Attribution for Conversational QA over Heterogeneous Data with RAG Systems

### v1

@@ 7.4. In-depth RAG analysis
In Table 4, we report several drill-down experiments with RAGonite configurations. Whenever one configuration choice was altered, the remaining values were held constant as per our default configuration specified in Sec. 7.2. The study leading to this table was also used to select our default configuration: whenever the best configuration following Precision@1 and Answer Relevance turn out to be different, we prefer the latter as the final response is more important to an end user. Specific observations below:
@@
    [Table linearization] In Rows $10-12$, verbalization of table records indeed makes table contents more digestible by an LLM: it is the winner when it comes to answer quality, but we also prefer this mode in the RAGonite interface owing to its easy scrutability by an end-user, over the plaintext content inside the table (in second place for both metrics).
@@
    [Generation LLM] In Rows $16-18$, we find that using GPT-4o led to superior results than Llama3.1 (we used the 8 billion parameter model [^5] https://huggingface.co/meta-llama/Llama-3.1-8B). But unfortunately the presence of large lists and table evidences in Confluence frequently led to the context window being exceeded for Llama, so its actual performance would be slightly better. The same LLM was used for both completed question and final answer generation.
@@
    [Embedding model] In Rows $19-21$, we observe that using bge-m3 embeddings [^6] https://huggingface.co/BAAI/bge-m3 for encoding text was better for our data than the recent text-embedding-3-small model [^7] https://openai.com/index/new-embedding-models-and-api-updates/ from OpenAI. The P@1 value for BGE in these rows ($0.427$) is lower than the global best in the other rows ($0.523$). This is because the BGE reranker was switched off for this experiment for fairness for both BGE and OpenAI embeddings, as the OpenAI embeddings were not compatible to be used with the BGE reranker. Notably, Answer Relevance is still the global best ($0.585$) since re-ranking only affects the relative ordering of evidences and hence alters P@1: the LLM GPT-4o still sees the same set of evidences and fortunately GPT-4o’s answering quality is unaffected by this relative order. As a sanity check, we can see that this P@1 of $0.427$ is the one for RRF in the Re-ranking experiment (Row 26). RRF is our fallback reranker for enabling hybrid search when the BGE re-ranker is switched off.
@@
    [Evidence ranking] In Rows $22-24$, hybrid search was found to be superior than lexical or dense search built inside ChromaDB. We used 10 evidences each from lexical and dense search to be fused for the hybrid variant. Hybrid search entails the re-ranking of evidences rerieved via lexical and dense search alone, as mentioned next.
@@
    [Evidence re-ranking] In Rows $25-27$, re-ranking evidences in the common pool of lexical and dense retrieval via the multilingual BGE re-ranking model bge-reranker-v2-m3888https://huggingface.co/BAAI/bge-reranker-v2-m3, to create the hybrid ranked list, was found to be better than a simple reciprocal rank fusion (RRF) approach where the final list was created by ordering via a new evidence score that is the sum of the reciprocals of the ranks of the evidence in the lexical and dense rankings.

### v2

@@ 7.4. In-depth RAG analysis
In Table 4, we report several drill-down experiments with RAGonite configurations. Whenever one configuration choice was altered, the remaining values were held constant as per our default configuration specified in Sec. 7.2. The study leading to this table was also used to select our default configuration. Specific observations are listed below:
@@
    [Table linearization] In Rows $10-12$, verbalization of table records indeed makes table contents more retrievable by a hybrid retriever (see below) as well as more digestible by an LLM, compared to natural linearization alternatives (the piped mode can be found in (Wang et al., 2024; Li et al., 2023a); markdown simply refers to the table in markdown format [^5] https://www.markdownguide.org/extended-syntax/#tables; HTML refers to the raw HTML of the table; plaintext is the remaining text of the table after stripping all HTML tags and attributes). But we also prefer verbalizations in the RAGonite interface owing to its easy scrutability by an end-user, over the HTML content of the table.
@@
    [Generation LLM] In Rows $16-18$, we find that using GPT-4o led to superior answer quality than Llama3.1 (we used the 8 billion parameter model [^6] https://huggingface.co/meta-llama/Llama-3.1-8B). But unfortunately the presence of large lists and table evidences in Confluence frequently led to the context window being exceeded for Llama, so its actual performance would be slightly better. The same LLM was used for both completed question and final answer generation. The P@1 is also lower for Llama ($0.480$ vs. $0.528$ for GPT-4o) because the retrieval in this case is carried out via the intent-explicit question generated by Llama, this is slightly inferior than the corresponding completions from GPT-4o. As requested in the prompt, the out-of-scope (OOS) message “The desired information cannot be found in the retrieved pool of evidence.” was triggered $199/600$ times ($33.2\%$) for the default RAGonite configuration with GPT-4o. Computing that the Precision@10 was $60.7\%$, an message would be expected $39.3\%$ of the times. Since this is very close to the actual measured rate of the OOS message, we can say that GPT-4o showed high instruction-following capabilities and did not use its parametric knowledge to generate its answers (undesirable in RAG).
@@
    [Embedding model] In Rows $19-21$, we observe that using bge-m3 embeddings for encoding text led to marginally better retrieval (P@1$=0.528$) for our data than the recent text-embedding-3-small model from OpenAI (P@1$=0.525$). While using the OpenAI model led to slightly better answer quality, better retrieval is the primary purpose of embeddings, leading to our choice of BGE as default. Moreover, BGE is used via local deployment, that we find preferable to an API for text-embedding-3-small because of latency and cost issues.
@@
    [Evidence ranking] In Rows $22-24$, hybrid search was found to be superior than lexical or dense search built inside ChromaDB. We used 10 evidences each from lexical and dense search to be fused for the hybrid variant. Hybrid search entails a re-ranking of evidences, as mentioned next.
@@
    [Evidence reranking] In Rows $25-27$, reranking evidences in the common pool of lexical and dense retrieval via the multilingual BGE re-ranking model to create the hybrid ranked list, was found to be better than a simple reciprocal rank fusion (RRF) approach. In RRF, the final list is created by ordering via a new evidence score that is the sum of the reciprocals of the ranks of the evidence in the lexical and dense rankings (a parameter $k=60$ is added to the original rank (Cormack et al., 2009) before computing the reciprocal). Note that RRF is always applied by default on the lexical and dense results to create the hybrid list: when the BGE reranker is applied, it works over the baseline list with RRF.